If it takes 5 mins to fill a 15 litre bucket from a tap of diameter 2/✓πcm then the reynolds number of the flow is (density of water is 1000kg/m3 and viscosity of water is 1000 pa.s

Grade:12

1 Answers

Vikas TU
6 years ago
We know that Reynolds number
R = ρvd/η
The volume of the water flowing out per second is
Q = v × πd2/4
v = 4Q/d2π
R = ρvd/η
R = 4ρQ/πdη
ρ = Density of water = 103 kg/m3
Q = 3 L / min = 3 × 10-3 / 60 = 5 × 10-5 m3/s
d = 2/√π × 10-2 m = 1.125 × 10-2 m
η = Viscosity = 10-3 Pa.s
So, R = (4 × 103 × 5 × 10-5)/(3.14 × 1.125 × 10-2 × 10-3)
R ≅ 5500